OpenStudy (anonymous):

y=4/3x+4 find slope and y-intercept

jimthompson5910 (jim_thompson5910):

The equation is in the form y = mx+b where m is the slope and (0,b) is the y-intercept. In this case, m = 4/3 and b = 4 So the slope is 4/3 and the y-intercept is (0,4)

OpenStudy (anonymous):

okay ty!!!!Again ;) if I wanted to graph this what might another y-intercept be?

OpenStudy (anonymous):

that funcion is linear. there should be only 1

OpenStudy (anonymous):

could it be none?/

jimthompson5910 (jim_thompson5910):

All lines only have ONE y-intercept. This is ALWAYS the case with any line. In fact, all functions only have one y-intercept. If you wanted another point on the other hand, plug in any x value to find its corresponding y value

OpenStudy (anonymous):

oh ty :)

jimthompson5910 (jim_thompson5910):

So for instance, if x = 3, then y = 4/3x + 4 y = (4/3)(3)+4 y = 12/3 + 4 y = 4+4 y = 8 So when x = 3, y = 8 So we have another point (3,8)

jimthompson5910 (jim_thompson5910):

So you now have two points (0,4) and (3,8)

OpenStudy (anonymous):

got it, thanks for the example :)